Solve the system by addition.
4x + 8y = –1
8x – 4y = –7

Answer by checkley75(3666)

8X-4Y=-7 NOW MULTIPLY BY -2 & ADD
-8X-16Y=2
--------------------------------
-20Y=-5
Y=-5/-20
Y=1/4 ANSWER
4X+8*1/4=-1
4X+2=-1
4X=-1-2
4X=-3
X=-3/4 ANSWER
PROOF
8*-3/4-4*1/4=-7
-24/4-1=-7
-6-1=-7
-7=-7